Dimeback Co. has total assets of $9,900,000 and a total asset turnover of 2.49 times. Assume the return on assets is 11 percent.

What is the company's sales? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Sales            $

What is the company's net income? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

Net income            $

What is its profit margin? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Profit margin             %
Answer #1

Total asset turnover=Sales/Total asset

Sales=(9,900,000*2.49)

=$24651000

ROA=Net income/Total assets

Net income=9,900,000*11%

=$1089000

Profit margin=Net income/Sales

=1089000/24651000

=4.42%(Approx).